Summary for evening shift 7/12/99

Renata Zukanovich-Funchal (zukanov@fnal.gov)
Tue, 13 Jul 1999 00:02:48 -0500

Shift Summary for evening shift, 07/12/99

On Shift: V. Prasad, C. Escobar, R. Funchal

Hours of beam available at all intensities: 8.4
Hours of beam available at nominal intensity: 8.4
Hours of E832 data (special runs not included): 7.4
Hours of special runs (laser scans, muon runs...): 0.2

Accounting of lost time: Tape changes, Epicure Hell ...
========================================================================
Runs ended this shift:

Run 13867 - BEAM_832 (24. M events)
Run 13868 - BEAM_MUON_4E12 (500k events)
Run 13869 - BEAM_832 (2. M events)

Tapes brought to FCC this shift: QK*125-127, QKB043
========================================================================
Accesses during shift: none
========================================================================
Overview of shift: 832 data taking at ~6.2 E12
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13867 15:30 Took over smooth running.

18:30 Dave N. changed 3 bottles Ar-Eth and bottle of CO2

21:12 Run ended, tape full. Tapes changed.
QK*125-127 and QKB043 taken out.
New tapes in place: QK*128-130 and QKB044

21:35 Discovered that mu run/laser scan destination logic
was wrong. S.T. fixed.

13868 22:03 New mu-run started.

22:13 Mu-run stopped. Switch to Magnet polarity = Neg.

13869 22:30 New BEAM_832 starts.

22:38 Booster problems, no beam.

22:45 Beam returns.

22:48 Ktev2 disk full. Onsplit requests QMU09 to be
mounted. We do so.

23:30 Run stopped. Problems with ktev1.

13870 23:35 New BEAM_832 starts.

========================================================================
Follow up Notes on Previous Shift
========================================================================
Items to Follow up on Next Shift

Do a laser scan, if beam is down.